OpenStudy (anonymous):

Acceleration (a) of a moving object is directly proportional to its velocity (v). If v is 4 when a is 10, find constant of variation.

OpenStudy (jdoe0001):

something is "directly proportional" to something else something = some ratio/proportion of something else

OpenStudy (anonymous):

????????

OpenStudy (jdoe0001):

accelaration = \(\bf Konstant \times \) velocity so we know that when v = 4, a = 10, thus accelaration = \(\bf Konstant \times \) velocity 10 = \(\bf Konstant \times \) 4 to find the "Konstant", just solve for that :)

OpenStudy (jdoe0001):

acceleration rather, anyhow hehe

OpenStudy (anonymous):

im sorry but im lost

OpenStudy (jdoe0001):

ok, what confused you?

OpenStudy (anonymous):

do i divide by 4 for the konsant

OpenStudy (anonymous):

konstant

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

now what do i do with the constant

OpenStudy (anonymous):

i got 2.5

OpenStudy (fifciol):

correct :)

OpenStudy (anonymous):

kk

OpenStudy (jdoe0001):

yes